# Global Climate Model Bias Correction Using Deep Learning

### v1

@@ Abstract
Climate change affects ocean temperature, salinity and sea level, which in turn impacts monsoons and ocean productivity. Future projections by Global Climate Models (GCM) based on shared socioeconomic pathways (SSP) from the Coupled Model Intercomparison Project (CMIP) are widely used to understand the effects of climate change. However, CMIP models have significant bias compared to reanalysis in the Bay of Bengal for the time period when both projections and reanalysis are available. For example, there is a 1.5C root mean square error (RMSE) in the sea surface temperature (SST) projections of the climate model CNRM-CM6 compared to the Ocean Reanalysis System (ORAS5). We develop a suite of data-driven deep learning models for bias correction of climate model projections and apply it to correct SST projections of the Bay of Bengal. We propose the use of three different deep neural network architectures: convolutional encoder-decoder UNet, Bidirectional LSTM and ConvLSTM. We also use a baseline linear regression model and the Equi-Distant Cumulative Density Function bias correction method for comparison and evaluating the impact of the new deep learning models. All bias correction models are trained using pairs of monthly CMIP6 projections and the corresponding month’s ORAS5 as input and output. Historical data (1950-2014) and future projection data (2015-2020) of CNRM-CM6 are used for training and validation, including hyperparameter tuning. Testing is performed on future projection data from 2021 to 2024. Detailed analysis of the three deep neural models has been completed. We found that the UNet architecture that is trained using a climatology-removed CNRM-CM6 projection as input and climatology-removed ORAS5 as output gives the best bias-corrected projections. Our novel deep learning-based method for correcting CNRM-CM6 data has a 15% reduction in RMSE compared to the widely used statistical correction technique, the Equidistant Cumulative Distribution Function (EDCDF).
@@ 1 Introduction
Several bias correction methods have been proposed in the literature for correcting GCM projections. The Equi-Distant Cumulative Distribution Function (EDCDF) method, a quantile-based mapping method for bias correction (Piani et al., 2010), is one of the most popular statistical bias correction methods used for the CMIP6 climate models. However, this method is limited to addressing errors in the simulated frequency distributions in each cell of the grid individually and has other issues such as instability and overfitting (Li et al., 2010; François et al., 2020).
@@
The remainder of the paper is organized as follows. Section 2 describes the mathematical problem statement, describes the literature survey of bias correction methods, discusses the choice of neural architectures, and describes the data sets. The main UNet model is developed in Section 3 and the other architecture choices including linear models and statistical baselines are provided in Section 4. Section 5 document the comparison of difference bias corrections methods. Finally, a summary and future directions are given in Section 6.
@@ 2.2 Related Work
Bias correction of GCM projections could be classified into three types: (i) interpolation methods, (ii) dynamical downscaling, and (iii) machine learning methods.
Interpolation methods
Interpolation is a widely used bias correction method that adjusts mean and variance using projection and reanalysis (or observational) data. Some of the main interpolation-based statistical methods are as follows. The delta change approach uses observations as a basis that produces future time series with dynamics similar to current conditions (Graham et al., 2007). The linear scaling approach adjusts the monthly mean values and offers corrected data with a consistent variability with the observation data (Graham et al., 2007). Local Intensity Scaling (LOCI) combines the advantages of linear scaling with a correction of wet day frequencies (precipitation threshold) (Schmidli et al., 2006). Although the linear scaling above accounts for a bias in the mean, it does not allow differences in the variance to be corrected. The power transformation is a nonlinear correction method in an exponential form to correct/adjust variance statistics (Leander and Buishand, 2007). A similar exponential variance scaling method adjusts both the variance and the mean in a stepwise manner (Chen et al., 2011). Another method called distribution mapping corrects the distribution function of the model value so that it matches the data (Sennikovs and Bethers, 2009). The most widely used is the quantile mapping technique (Piani et al., 2010) that adjusts the distributional properties of the model more closely to match those of the historical observations. There are different versions of quantile mapping methods available, of which the Equi-Distant Cumulative Distribution Function (EDCDF) method is the most popular for bias correction of CMIP6 projections (Li et al., 2010). Some studies have shown that EDCDF effectively reduces biases in GCM projections Sachindra et al. (2012); Yang et al. (2018); Mondal et al. (2021). In this work, we have used EDCDF as the baseline for evaluating our neural model to correct the CMIP6 projections (Section 4.3).
@@
Machine learning (ML) algorithms can make data-driven predictions using supervised learning that involves learning a mapping function between input and output variables. With the growing popularity and advances in machine learning, there has been an increase in the use of ML techniques with environmental data (Reichstein et al., 2019). These techniques are used to analyze and model complex ocean and atmospheric data, such as temperature, precipitation, salinity, and currents, to improve our understanding of the atmosphere and ocean dynamics and their impact on climate and weather patterns (Ahijevych et al., 2016; Boukabara et al., 2020). Furthermore, ML is being used to develop more accurate and efficient model forecasting, which can have important implications for various sectors such as agriculture, energy, transportation, infrastructure planning, fishing, and offshore energy production (Reichstein et al., 2019). Deep neural networks, specifically convolutional and transformer models, have the potential to tackle spatio-temporal problems (Labe et al., 2024; Baño-Medina et al., 2024). Recently, deep learning has been used to correct for bias in Tmin, Tmax (Wang and Tian, 2022), precipitation (Fulton et al., 2023), and wind energy (Zhang and Li, 2021). As far as we are aware, bias correction in CMIP6 has not been attempted using deep learning methods for the Bay of Bengal region. From the results cited above in other areas of Earth system modeling, the scope for machine learning for CMIP6 bias correction appears to be high. In this paper, we develop a suite of deep neural networks to correct errors in CMIP6 projections in the Bay of Bengal and identify the best architecture for this purpose.
@@ 2.5 Training, Validation and Testing
The monthly climatology is derived from the ORAS5 reanalysis and subtracted from the input and target data. Our ablation study indicates that this anomaly correction produces optimal results. All monthly climatology-removed data from CNRM-CM6 and ORAS5 were resized to $128\times 128$ for our UNet model. This choice is to ensure that the dimensions have a power of 2 and that the original data resolution is not lost. ORAS5 is used as the target for all SSP scenarios.
@@ 3 UNet: Fully Convolutional Encoder-Decoder Architecture
The UNet architecture takes as input a $128\times 128$ time slice of the monthly projection by CNRM-CM6. The CNRM-CM6 and ORAS5 data are available with a resolution of 85$\times$85 in the Bay of Bengal study domain. These fields are bilinearly interpolated to $128\times 128$. The processing is done by using a fully convolutional encoder-decoder structure. The encoder path comprises five downsampling stages using $3\times 3$ convolutional layers with tanh activation and dropout, followed by $2\times 2$ max pooling, with the number of filters doubling at each stage (32, 64, 128, 256, 512). The bottleneck contains two $3\times 3$ convolutions with 1024 filters. The decoder mirrors this structure with five upsampling stages using $2\times 2$ transposed convolutions, where each stage combines the upsampled features with the corresponding encoder features through skip connections (concatenate operation). Each decoder block includes two $3\times 3$ transpose convolutional layers with tanh activation, and the final output is produced by a $1\times 1$ convolution. The exact choice of the number of filters at each stage was reached during hyperparameter tuning on the validation data. We implement Python-based data processing tools and a UNet neural network implementation for bias correction of CNRM-CM6 SST and make our code available on GitHub111 https://github.com/AbhiPasula/Climate-Model-SST-Bias-Correction-Toolkit.
@@ 3.1 Data Pre-processing
Through extensive experimentation we found out that pre-processing the input climate projections and output ORAS5 reanalysis using the monthly climatology from ORAS5 achieves the best bias correction results. Using ORAS5 data from 1958 to 2014, we calculate the mean of each month. Then, this calculated mean is removed from the climate model projections for each month and the corresponding ORAS5 target. During inference, the monthly climatology is added back to the output of our UNet model. We hypothesize that using climatology-removed data for training and inference allows the neural model to better learn the distribution of data.
@@ 4 Other Architecture Choices and Baselines
To arrive at the final UNet model, we performed several ablation studies and also experimented with other architectural choices. We trained two variants of sequence-to-sequence deep learning models based on recurrent neural architectures. First, we flattened the ocean grid points in a row-major (latitude) and column-major (longitude) form and used a bidirectional LSTM model. In this choice, we used the projection of one month as input, and the corrected projection of the same month as the output was obtained. In another sequence-to-sequence model, we used a ConvLSTM spatio-temporal model that accepted a sequence of 4 month projections as input and produced a 1 month corrected projection as the output. For baseline, we developed a linear regression model that operates on a flattened vector of ocean grid points. We also used the statistical EDCDF method as a baseline to compare the performance of all machine learning models.
@@ 5.1 Comparison of models in test period
The RMSE of SST projections in the Bay of Bengal for the test years is shown in Table 1. The raw CNRM-CM6 climate model shows the highest RMSE values (1.22-1.34). The RMSE comparison across different GCM correction methods reveals UNet’s superior performance in all tested scenarios. UNet consistently achieves the lowest RMSE values, ranging from 0.4920 for SSP5 to 0.5018 for SSP1, indicating more accurate temperature predictions than other methods. Linear regression shows considerably higher RMSE values (0.77-0.82) than EDCDF (0.62-0.79), highlighting the limitations of simple statistical approaches for complex climate patterns. BiLSTM and ConvLSTM demonstrate comparable error metrics (around 0.53-0.59), but fall short of UNet’s accuracy. These quantitative results provide strong evidence that UNet offers the most reliable approach to correcting climate model errors in the Bay of Bengal region.
5.2 Comparison of models in future projections
The long term annual sea surface temperature (SST) projections for the BoB from 1958 to 2100 for four different SSPs, SSP1 (SSP1-2.6), SSP2 (SSP2-4.5), SSP3 (SSP3-7.0), and SSP5 (SSP5-8.5), are shown in Fig. 4. Historical CNRM-CM6 (CMIP6) consistently underestimates the temperatures compared to ORAS5 reanalysis. The long-term trend of SST shows significant variations between different correction methods and SSP scenarios. SST projections indicate substantial warming, with the UNet-corrected SSP5 scenario projecting increases of 3-4C by 2100, while the more conservative SSP1 scenario suggests modest warming of 1-1.5C. Despite their varying intensities, the consistent upward trend across all scenarios strongly suggests continued ocean warming through the twentieth century, with the magnitude of warming heavily dependent on future greenhouse gas emissions and climate policy decisions. Comparison between SSP scenarios reveals distinct warming trajectories that reflect their underlying socioeconomic and emission assumptions. SSP5-8.5 demonstrates the most dramatic warming, with temperatures reaching approximately 33oC by 2100 (Fig. 4(d)), SSP3-7.0 follows a similar but slightly moderated pattern, with end-century temperatures around 32oC (Fig. 4(c)), SSP2-4.5 shows an intermediate warming path with temperatures approaching 31oC by 2100 (Fig. 4(b)), SSP1-2.6, representing stringent mitigation measures, shows the least warming with temperatures stabilizing around 30oC after 2060 (Fig. 4(a)). The magnitude of warming indicated by the UNet-corrected projections is less severe than that indicated by the EDCDF-corrected SST projections. The BiLSTM correction fails to effectively capture the variability in annual mean projections, producing overly smoothed temperature trends. ConvLSTM projections align more closely with EDCDF-corrected projections. UNet captures both the magnitude and variability of SST changes in the GCM bias correction, suggesting that future warming in the Bay of Bengal may differ from previous projections, with implications for regional climate patterns, monsoon behavior, and marine ecosystems. Overall, we see that the UNet correction has the lowest RMSE and captures the variability accurately.

### v2

@@ Abstract
Climate change affects ocean temperature, salinity, and sea level, which in turn impacts monsoons and ocean productivity. Future projections by Global Climate Models (GCM) based on shared socioeconomic pathways (SSP) from the Coupled Model Intercomparison Project (CMIP) are widely used to understand the effects of climate change. However, CMIP models have a significant bias compared to reanalysis in the Bay of Bengal. For example, there is a 1.5C root mean square error (RMSE) in the sea surface temperature (SST) projections of the climate model CNRM-CM6 compared to the Ocean Reanalysis System (ORAS5). We develop a suite of data-driven deep learning models for bias correction of climate model projections and apply it to correct SST projections of the Bay of Bengal. We propose the use of three different deep neural network architectures: a convolutional encoder-decoder UNet, a Bidirectional LSTM, and a ConvLSTM. We used a baseline linear regression model and the Equi-Distant Cumulative Density Function (EDCDF) bias correction method for comparison and evaluation of the impact of the new deep learning models. All bias correction models are trained using pairs of monthly CMIP6 projections and the corresponding month’s ORAS5 as input and output. Historical data (1950-2014) and future projection data (2015-2020) of CNRM-CM6 are used for training and validation, including hyperparameter tuning. Testing is performed on future projection data from 2021 to 2024. We found that the UNet architecture that is trained using a climatology-removed CNRM-CM6 projection as input and climatology-removed ORAS5 as output gives the best bias-corrected projections. Our novel deep learning-based method for correcting CNRM-CM6 data has a 15% reduction in RMSE and a 12% increase in the pattern correlation coefficient compared to EDCDF.
@@ 1 Introduction
Several statistical bias correction methods have been proposed in the literature for correcting GCM projections (Volpi et al., 2024). The Equi-Distant Cumulative Distribution Function (EDCDF) method, a quantile-based mapping method for bias correction (Piani et al., 2010), is one of the most popular statistical bias correction methods used for the CMIP6 climate models. However, this method is limited to addressing errors in the simulated frequency distributions in each cell of the grid individually and has other issues such as instability and overfitting (Li et al., 2010; François et al., 2020).
@@
The remainder of the paper is organized as follows. Section 2 describes the mathematical problem statement, describes the literature survey of bias correction methods, discusses the choice of neural architectures, and describes the data sets. The main UNet model is developed in Section 3, and the other architecture choices including linear models and statistical baselines are provided in Section 4. Section 5 documents the comparison of different bias correction methods. Finally, a summary and future directions are given in Section 6.
@@ 2.1 Mathematical Problem Statement
The above formulation is an operator learning framework for bias correction. We do not explicitly train the neural operator with a target to correct any specific moments such as mean, variance, or higher-order moments; rather the mean square error is minimized between the neural operator’s output and the reanalysis data. This approach forces the operator to learn the distributional properties (Lu et al., 2021; Kovachki et al., 2023; Raonic et al., 2023).
@@ 2.2 Related Work
Bias correction of GCM projections could be classified into three types: (i) statistical methods, (ii) dynamical downscaling, and (iii) machine learning methods.
Statistical methods
Statistical bias correction methods can be categorized as univariate and multivariate methods. Univariate methods address individual variables on their own, while multivariate methods handle biases for each variable and also modify the interdependencies among variables in the climate model (Piani et al., 2010; Piani and Haerter, 2012; Cannon et al., 2015; Maraun, 2016; Volpi et al., 2024). Interpolation is a widely used bias correction method that adjusts mean and variance using projection and reanalysis (or observational) data. Some extensively used univariate interpolation statistical methods are explained below. The delta change approach uses observations as a basis that produces future time series with dynamics similar to current conditions (Graham et al., 2007). The linear scaling approach adjusts the monthly mean values and offers corrected data with a consistent variability with the observation data (Graham et al., 2007). Local Intensity Scaling (LOCI) combines the advantages of linear scaling with a correction of wet day frequencies (precipitation threshold) (Schmidli et al., 2006). Although the linear scaling above accounts for a bias in the mean, it does not allow differences in the variance to be corrected. The power transformation is a nonlinear correction method in an exponential form to correct/adjust variance statistics (Leander and Buishand, 2007). A similar exponential variance scaling method adjusts both the variance and the mean in a stepwise manner (Chen et al., 2011). Another method called distribution mapping corrects the distribution function of the model value so that it matches the data (Sennikovs and Bethers, 2009). The most widely used is the quantile mapping technique (Piani et al., 2010) that adjusts the distributional properties of the model more closely to match those of the historical observations. There are different versions of quantile mapping methods available, of which the Equi-Distant Cumulative Distribution Function (EDCDF) method is the most popular for bias correction of CMIP6 projections (Li et al., 2010).
In multivariate methods, conditional binning, a two-dimensional conditional bias correction approach for temperature and precipitation, was proposed by (Piani and Haerter, 2012). The temperature is first corrected by using quantile mapping. Second, precipitation and temperature couples are assigned to one of several temperature bins. Finally, using quantile mapping, precipitation data inside each temperature bin are bias-adjusted. Another category of multivariate methods is iterative; the core concept is to repeatedly apply univariate quantile mapping and multivariate correction (Cannon, 2016) across multiple time scales (Mehrotra and Sharma, 2016).
In nested bias correction, the distributional and persistence bias is corrected from fine to progressively longer time scales (Johnson and Sharma, 2012). Frequency-based bias correction works in the frequency space and is independent of specific time scales (Nguyen et al., 2016). The wavelet-based signal processing bias correction approach was implemented for the GCM future time series and described the discontinuity in the trend from current to future climate projections (Kusumastuti et al., 2021). Studies have shown that quantile mapping, especially EDCDF, effectively reduces biases in GCM projections (Sachindra et al., 2012; Yang et al., 2018; Mondal et al., 2021). In this work, we have used EDCDF as the baseline for evaluating our neural model to correct the CMIP6 projections (Section 4.3).
@@
Machine learning (ML) algorithms can make data-driven predictions using supervised learning that involves learning a mapping function between input and output variables. With the growing popularity and advances in machine learning, there has been an increase in the use of ML techniques with environmental data (Reichstein et al., 2019). These techniques are used to analyze and model complex ocean and atmospheric data, such as temperature, precipitation, salinity, and currents, to improve our understanding of the atmosphere and ocean dynamics and their impact on climate and weather patterns (Ahijevych et al., 2016; Boukabara et al., 2020). Furthermore, ML is being used to develop more accurate and efficient model forecasting, which can have important implications for various sectors such as agriculture, energy, transportation, infrastructure planning, fishing, and offshore energy production (Reichstein et al., 2019). Deep neural networks, specifically convolutional and transformer models, have the potential to tackle spatio-temporal problems (Labe et al., 2024; Baño-Medina et al., 2024). Recently, deep learning has been used to correct for bias in Tmin, Tmax (Wang and Tian, 2022), precipitation (Fulton et al., 2023), and wind energy (Zhang and Li, 2021). In the Indian region, the near-surface temperature (Dutta and Bhattacharjya, 2022) and ISMR (Sharma et al., 2024) were bias corrected using deep learning models. To our knowledge, bias correction of the CMIP6 ocean forecasts for the Bay of Bengal region has not been attempted using deep learning methods From the results cited above in other areas of Earth system modeling, the scope for machine learning for CMIP6 bias correction appears to be high. In this paper, we develop a suite of deep neural networks to correct errors in CMIP6 projections in the Bay of Bengal and identify the best architecture for this purpose.
@@ 2.5 Training, Validation and Testing
The monthly climatology is derived from the ORAS5 reanalysis and subtracted from the input and target data. Our ablation study indicates that this climatology removal produces optimal results. Such pre-processing is widely used in the GCM bias correction methods (Tabor and Williams, 2010; Kim et al., 2014; Navarro-Racines et al., 2020; Xu et al., 2021). All monthly climatology-removed data from CNRM-CM6 and ORAS5 were resized to $128\times 128$ for our UNet model. This choice is to ensure that the dimensions have a power of 2 and that the original data resolution is not lost due to the symmetric encoder-decoder structure and skip connections of the UNet. ORAS5 is used as the target for all SSP scenarios.
@@ 3 UNet: Fully Convolutional Encoder-Decoder Architecture
The UNet architecture takes as input a $128\times 128$ time slice of the monthly projection by CNRM-CM6. The CNRM-CM6 and ORAS5 data are available with a resolution of $\times85$ in the Bay of Bengal study domain. These fields are bilinearly interpolated to $128\times 128$. The processing is done by using a fully convolutional encoder-decoder structure. The encoder path comprises five downsampling stages using $3\times 3$ convolutional layers with tanh activation and dropout, followed by $2\times 2$ max pooling, with the number of filters doubling at each stage (32, 64, 128, 256, 512). The bottleneck contains two $3\times 3$ convolutions with 1024 filters. The decoder mirrors this structure with five upsampling stages using $2\times 2$ transposed convolutions, where each stage combines the upsampled features with the corresponding encoder features through skip connections (concatenate operation). Each decoder block includes two $3\times 3$ transpose convolutional layers with tanh activation, and the final output is produced by a $1\times 1$ convolution. The exact choice of the number of filters at each stage was reached during hyperparameter tuning on the validation data. We implement Python-based data processing tools and a UNet neural network implementation for bias correction of CNRM-CM6 SST and make our code available on GitHub (https://github.com/AbhiPasula/Climate-Model-SST-Bias-Correction-Toolkit).
The above convolutional neural operator (Raonic et al., 2023) uses convolutional layers to account for spatial dependencies in bias correction, thereby enhancing the overall accuracy of the deep model.
@@ 3.1 Data Pre-processing
Through extensive experimentation, we found that preprocessing the input climate projections and output ORAS5 reanalysis using the monthly climatology from ORAS5 achieves the best bias correction results. Using ORAS5 data from 1958 to 2014, we calculate the mean of each month. Then, this calculated mean is removed from the climate model projections for each month and the corresponding ORAS5 target. During inference, the monthly climatology is added back to the output of our UNet model. We hypothesize that using climatology-removed data for training and inference allows the neural model to better learn the distribution of data.
@@ 4 Other Architecture Choices and Baselines
To arrive at the final UNet model, we performed several ablation studies and also experimented with other architectural choices. We trained two variants of sequence-to-sequence deep learning models based on recurrent neural architectures. First, we flattened the ocean grid points in a row-major (latitude) and column-major (longitude) form and used a bidirectional LSTM model. In this choice, we used the projection of one month as input, and the corrected projection of the same month as the output was obtained. In another sequence-to-sequence model, we used a ConvLSTM spatio-temporal model that accepted a sequence of 4 month projections as input and produced a 1 month corrected projection as the output. For baseline, we developed a linear regression model that operates on a flattened vector of ocean grid points. We also used the statistical EDCDF method as a baseline to compare the performance of all machine learning models. In the UNet ablation study, four UNet architectural variants were systematically evaluated to assess the contribution of different design components to model performance. The supporting information provides a comprehensive description of further ablation studies for the UNet architecture. Section S2 contains a thorough analysis of the UNet ablation study, while Section S7 presents findings from an additional ablation study that uses HadGem3 data as input to the UNet model trained on CNRM-CM6 data.
@@ 5.1 Comparison of models in test period
The RMSE and pattern correlation coefficient (PCC) of SST projections in the Bay of Bengal for the test years are shown in Table 1 and Table 2 respectively. The raw CNRM-CM6 climate model shows the highest RMSE values (1.22-1.34) and lowest PCC values (0.57-0.58), indicating significant bias and weak correlation with ORAS5 reanalysis. Among the correction methods, UNet consistently shows superior performance across both metrics, achieving the lowest RMSE values ranging from 0.4920 for SSP5 to 0.5018 for SSP1, while simultaneously maintaining the highest PCC values (0.7613-0.7768), demonstrating both superior accuracy and strong correlation with ORAS5. Linear regression shows the weakest performance among correction methods with considerably higher RMSE values (0.77-0.82) and lower PCC values (0.6739-0.6990), highlighting the limitations of simple statistical approaches for capturing complex climate patterns. EDCDF achieves moderate performance with RMSE values of 0.62-0.79 and PCC values of 0.6515-0.6871, while BiLSTM and ConvLSTM demonstrate comparable but inferior performance to UNet, with RMSE values around 0.53-0.59 and PCC values of 0.6817-0.7151. These quantitative results provide strong evidence that UNet offers the most reliable approach for correcting climate model errors in the Bay of Bengal region, combining both the highest accuracy (lowest RMSE) and strongest correlation (highest PCC) with observed SST.
@@
Table 2: PCC in SST projections of test years (2021-2024) obtained by different CNRM-CM6 correction methods. The best values are highlighted in bold text.
Scenario CNRM
Linear
Regression
EDCDF UNet ConvLSTM BiLSTM
SSP1 0.5747 0.6990 0.6812 0.7768 0.6845 0.6941
SSP2 0.5721 0.6897 0.6750 0.7613 0.7151 0.7049
SSP3 0.5743 0.6866 0.6809 0.7615 0.7058 0.6952
SSP5 0.7285 0.6739 0.6515 0.7666 0.6992 0.6817
5.2 Comparison of models in future projections
The long term annual sea surface temperature (SST) projections for the Bay of Bengal from 1958 to 2100 for four different SSPs, SSP1 (SSP1-2.6), SSP2 (SSP2-4.5), SSP3 (SSP3-7.0), and SSP5 (SSP5-8.5), are shown in Fig. 4. Historical CNRM-CM6 (CMIP6) consistently underestimates the temperatures compared to ORAS5 reanalysis. A noticeable break in trend is evident when comparing historical to future climate projections. This artifact is expected for statistically bias-corrected time series projections (Kusumastuti et al., 2021). The long-term trend of SST shows significant variations between different correction methods and SSP scenarios. SST projections indicate substantial warming, with the UNet-corrected SSP5 scenario projecting increases of 3-4C by 2100, while the more conservative SSP1 scenario suggests modest warming of 1-1.5C. Despite their varying intensities, the consistent upward trend across all scenarios strongly suggests continued ocean warming through the twentieth century, with the magnitude of warming heavily dependent on future greenhouse gas emissions and climate policy decisions. Comparison between SSP scenarios reveals distinct warming trajectories that reflect their underlying socioeconomic and emission assumptions. SSP5-8.5 demonstrates the most dramatic warming, with temperatures reaching approximately 33oC by 2100 (Fig. 4(d)), SSP3-7.0 follows a similar but slightly moderated pattern, with end-century temperatures around 32oC (Fig. 4(c)), SSP2-4.5 shows an intermediate warming path with temperatures approaching 31oC by 2100 (Fig. 4(b)), SSP1-2.6, representing stringent mitigation measures, shows the least warming with temperatures stabilizing around 30oC after 2060 (Fig. 4(a)). The magnitude of warming indicated by the UNet-corrected projections is less severe than that indicated by the EDCDF-corrected SST projections. The EDCDF correction method, as it is based on quantiles, often inflates trends, resulting in significant deviation in the long-term SST trend in the Bay of Bengal (Cannon et al., 2015; Maraun, 2016). The BiLSTM correction does not effectively capture the variability in the annual mean projections, producing overly smoothed temperature trends. ConvLSTM projections align more closely with EDCDF-corrected projections. UNet captures both the magnitude and variability of SST changes in the GCM bias correction, suggesting that future warming in the Bay of Bengal may differ from previous projections, with implications for regional climate patterns, monsoon behavior, and marine ecosystems. Overall, we see that the UNet correction has the lowest RMSE and captures the variability accurately.
Our comparison results show that for the task of bias correction, the UNet architecture, a simple convolutional neural operator (Raonic et al., 2023), is ideal. Other sequential processing models such as ConvLSTM and the Bi-LSTM models are sub-optimal for this application. In this study, we restricted our focus to convolutional and recurrent architectures and did not use transformer models (Liu et al., 2021; Khan et al., 2022; Han et al., 2022). This choice is primarily due to computational demands and limited data availability for training and validation, which are insufficient to train large transformer models. For example, the Swin transformer has a much higher parameter count than our UNet (Xiang et al., 2022; Zhong et al., 2024). However, in the future, these heavy models could be implemented for the bias correction problem and compared against the convolutional neural operator used in this paper.